OpenStudy (anonymous):

Use the quadratic formula to solve 4x2 + x - 1 = 0

hartnn (hartnn):

Compare your quadratic equation with \(ax^2+bx+c=0\) find \[a=...?\\b=...?\\c=...?\\\] \[\\~ \\ \sqrt{b^2-4ac}=...?\] then the two roots of x are: \(\huge{x_{1,2}=\frac{-b \pm \sqrt{b^2-4ac}}{2a}}\)

OpenStudy (johnweldon1993):

\[-b \pm \sqrt{b ^{2} - 4ac}/2a\]

OpenStudy (anonymous):

A. {.0.5, 0.7} B. {-0.3, -0.1} C. {0.2, 0.4} D. {-0.6, 0.4} This what i have to pick from

hartnn (hartnn):

yes, for that u need to use the formula, so, what values of a,b,c you got ?

OpenStudy (anonymous):

a.4x^2 b.x c. 1 ???

hartnn (hartnn):

a,b,c are constants so, a= 4 b=1 and c=-1 got this ?

OpenStudy (anonymous):

Oh okay. Yes i got it . Thank you

hartnn (hartnn):

welcome :) ask if you need further help with this.